OpenStudy (maddy1251):

Use the rules of logarithms to expand log[(x+3)^4(x-5)^7]

OpenStudy (owlcoffee):

Like this? \[\log[(x+3)^{4(x-5)^7}]\]

OpenStudy (maddy1251):

\[LOG[(x+3)^{4} (x-5)^{7}]\]

OpenStudy (maddy1251):

I should of posted my question and then wrote that out :P

OpenStudy (welshfella):

use the following 2 rules log ab = log a + log b and log a^n = n log a

OpenStudy (maddy1251):

@welshfella are you saying this\[\log (ab)=\log (x+3)^{4} +\log(x-5)^{7}\]

OpenStudy (owlcoffee):

okay, that makes things easier. \[\log[(x+3)^4(x-5)^7]\] You see, most of the times we want to expand a logarithmic math expression we use the properties of the logarithm. In this case we will use a logarithmic property that involves the product of the internal logarithmands. \[\log_{n} AB=\log_{n} A + \log_{n}B \] Well, in the case of this excercise we can apply it, since the base is 10: \[\log[(x+3)^4(x-5)^7]\] \[\iff \log (x+3)^4+\log(x-5)^7\] This of course is not ready yet, because we still have those exponents, from that we will use the property: \[\log_{n} A^m = m(\log_{n} A)\] So, we can take down those exponents by using this property, which comes in quite useful when dealing with logarithms: \[ \log (x+3)^4+\log(x-5)^7\] \[\iff 4\log(x+3)+7\log(x-5)\]

OpenStudy (welshfella):

you are on the way to answering it now you apply the second rrule

OpenStudy (welshfella):

OwlCoffee has it

OpenStudy (maddy1251):

@welshfella Ah you started and then @Owlcoffee finished. I saw where you were going, it was a matter of taking the exponents after you got it into the first form! So simple..

OpenStudy (maddy1251):

@Owlcoffee I will medal you, can you medal @welshfella

OpenStudy (welshfella):

yes Not difficult

OpenStudy (maddy1251):

@welshfella Thank you!! :)

OpenStudy (welshfella):

yw

OpenStudy (ninjaslice):

I put in the right answer though with my image right?